Find the distance between the following perpendicular lines x+3y=7 and (sqr(1+ sqr64)x +5y +5=0


1
Expert's answer
2021-09-09T00:29:17-0400

The lines "x+3y=7" and "(\sqrt{1+\sqrt{64}})x+5y+5=0" are neither perpendicular nor parallel lines


"y=-\dfrac{1}{3}x+\dfrac{7}{3}, slope_1=-\dfrac{1}{3}"

"y=-\dfrac{\sqrt{1+\sqrt{64}}}{5}x-1, slope_2=-\dfrac{9}{5}"

"slope_1\not=slope_2"

"slope_1(slope_2)=-\dfrac{1}{3}(-\dfrac{9}{5})=\dfrac{3}{5}\not=-1"

The lines "x+3y=7" and "(\sqrt{1+\sqrt{64}})x+5y+5=0" intersect.

We the shortest distance between the given lines is zero.


When two parallel lines ("slope_1=slope_2" ) are given by


"ax+by+c_1=0"

"ax+by+c_2=0"

the distance between them can be expressed as


"d=\dfrac{|c_2-c_1|}{\sqrt{a^2+b^2}}"